# Route to Chaos and Unified Dynamical Framework of Multi-Species Ecosystems

### v1

@@ Abstract
Conclusions. It has long been known that multi-species ecosystems described by sets of coupled ordinary differential equations of the type given in Eq. (1) may exhibit any dynamical behavior Sma76 . Here we have emphasized a route joining the previously observed phases governed by attractive fixed-points to those exhibiting persistent dynamics. Because species in ecosystems and food webs interact with one another in a necessarily asymmetric way, fixed-point instabilities may occur via a Hopf bifurcation. The resulting limit cycles preserve biodiversity, since at least for some range of interaction variability, their stability does not necessitate species extinctions.
@@
Models of theoretical ecology are not expected to precisely reflect ecosystem behaviors. Instead, they may shed light on different, observed behaviors of real or lab-controlled ecosystems at a statistical level. In that respect, our work emphasizes the emergence of chaotic behaviors of species population in ecosystems. Recent analysis of population time series found evidence of chaotic behavior in at least 30% of the studied populations Rog21 . Fig. 2 further emphasizes parametrically sizable regimes with small Lyapunov exponents, which has also been observed  Rog21 ; Ell95 . Other regimes with strong sensitivity to even minor parametric changes [E.g. at $\sigma\gtrsim 3$ in Fig. 2(f)] are characteristic of systems close to criticality as discussed in Refs. Sol02 ; Mor11 . Since criticality slows down the dynamics, the behavior of ecosystems in this latter regime is transient by nature and not governed by any long-time asymptotic. That ecosystems are governed by transient behaviors has been postulated in Refs. Hu22 ; Has94 ; Schr03 ; Mor16 ; Has18 . The emergence of Hopf bifurcations at negative values of $\gamma$ finally explains recent results where ecosystems with dominating predator-prey interactions display oscillatory behaviors in their population dynamics Mam22 . The present work illustrates that all the different, observed or theoretically postulated behaviors naturally emerge from a single, unified model, without the need for exogeneous intervention. The task at hand now is obviously to try and determine, at least qualitatively, model parameters corresponding to specifically observed ecosystems. Our work has simplified that task in that it identified the two key model parameters $\sigma$ and $\gamma$ driving transitions between different dynamical behaviors.
@@
There are several important direction in which our work should be extended. Among them, we mention first, that we are currently investigating the frequency of occurence of strange attractors vs. restabilization after species extinction at large $\sigma$. Second, further investigations should extend our results to interaction matrices reflecting more realistic topologies of known ecological networks Fri17 . Third, investigating species distributions may give precious information on the conditions under which biodiversity and rarity may coexist Sol02 . Fourth, ecosystem parameters are modified by climate changes Ral10 ; Eng11 ; Par19 . Investigating changes in ecosystem functioning under climate-induced changes in trophic interactions is of paramount interest. It would evidently have a strong influence on ecosystem functioning in the critical and chaotic regimes with strong parameter sensitivity. There are certainly many other interesting extensions.
@@
Acknowledgments. We thank Christian Mazza and Xavier Richard for introducing us to theoretical ecology and population dynamics. RD was supported by the Swiss National Science Foundation, under grant nr. 200021_215336.

### v2

@@ IV Conclusions
It has long been known that multi-species ecosystems described by sets of coupled ordinary differential equations of the type given in Eq. (1) may exhibit any dynamical behavior. Smale (1976) Here we have emphasized a route joining the previously observed phases governed by attractive fixed-points to those exhibiting persistent dynamics. Because species in ecosystems and food webs interact with one another in a necessarily asymmetric way, fixed-point instabilities may occur via a Hopf bifurcation. The resulting limit cycles preserve biodiversity, since at least for some range of interaction variability, their stability does not necessitate species extinctions.
@@
Models of theoretical ecology are not expected to precisely reflect ecosystem behaviors. Instead, they may shed qualitative light on different, observed behaviors of real or lab-controlled ecosystems at a statistical level. In that respect, our work emphasizes the existence of various dynamical behaviors, depending on the two parameters $\sigma$ and $\gamma$ governing interspecies interactions. Recent analysis of population time series found evidence of chaotic behavior in at least 30% of the studied populations. Rogers, Johnson, and Munch (2022) Fig. 2 further emphasizes parametrically sizable regimes with small Lyapunov exponents, which has also been observed. Rogers, Johnson, and Munch (2022); Ellner and Turchin (1995) Other regimes with strong sensitivity to even minor parametric changes [e.g., at $\sigma\gtrsim 3$ in Fig. 2(f)] are characteristic of systems close to criticality as discussed in Refs. Solé et al., 2002; Mora and Bialek, 2011. Since criticality slows down the dynamics, the behavior of ecosystems in this latter regime is transient by nature and not governed by any long-time asymptotic. That ecosystems are governed by transient behaviors has been postulated in Refs. Hu et al., 2022; Hastings and Higgins, 1994; Schreiber, 2003; Morozov, Banerjee, and Petrovskii, 2016; Hastings et al., 2018. The emergence of Hopf bifurcations at negative values of $\gamma$ finally explains recent results where ecosystems with dominating predator-prey interactions display oscillatory behaviors in their population dynamics. Mambuca, Cammarota, and Neri (2022)
The present work illustrates that all the different, observed or theoretically postulated behaviors naturally emerge from a single, unified model, without the need for exogeneous intervention. The task at hand now is obviously to try and determine, at least qualitatively, model parameters corresponding to specifically observed ecosystems. Our work has simplified that task in that it identified the two key model parameters $\sigma$ and $\gamma$ driving transitions between different dynamical behaviors.
@@
There are several important direction in which our work should be extended. Among them, we mention first, that we are currently investigating the frequency of occurence of strange attractors vs. restabilization after species extinction at large $\sigma$. Second, further investigations should extend our results to interaction matrices reflecting more realistic topologies of known ecological networks. Fried, Shnerb, and Kessler (2017) Third, investigating species distributions may give precious information on the conditions under which biodiversity and rarity may coexist. Solé et al. (2002) Fourth, ecosystem parameters are modified by climate changes. Rall et al. (2010); Englund et al. (2011); Parain et al. (2019) Investigating changes in ecosystem functioning under climate-induced changes in trophic interactions is of paramount interest. It would evidently have a strong influence on ecosystem functioning in the critical and chaotic regimes with strong parameter sensitivity. There are certainly many other interesting extensions.
Supplementary Material
Section I of the supplementary material provides a numerical estimate of the probability that the extreme eigenvalue of the matrix in Eq. (2) is complex. In Sec. II of the supplementary material, we numerically illustrate various dynamical behaviors that are observed in the system of Eq. (1).
Acknowledgments
@@
We thank Christian Mazza and Xavier Richard for introducing us to theoretical ecology and population dynamics. RD was supported by the Swiss National Science Foundation, under grant nr. 200021_215336.
Author declaration
Conflict of interest
The authors have no conflict to disclose.
Author contributions
Robin Delabays: Conceptualization (support); Formal analysis (support); Investigation (support); Methodology (support); Visualization (equal); Writing – original draft (support); Writing – review & editing (equal).
Philippe Jacquod: Conceptualization (lead); Formal analysis (lead); Investigation (lead); Methodology (lead); Visualization (equal); Writing – original draft (lead); Writing – review & editing (equal).
Data availability statement
The data that support the findings of this study are available within the article.
Supplementary material
IV.1 Extreme eigenvalue of a real asymmetric random matrix
The occurence of a Hopf bifurcation requires that the first eigenvalue to cross the imaginary axis has a finite imaginary part – in which case one actually has a pair of different, complex-conjugated eigenvalue. Therefore, the probability of a Hopf bifurcation is larger, if the probability that the extreme eigenvalue of the stability matrix has a finite imaginary part is larger. Fig. 5 shows this probability for the matrix $\mathbb{A}$ of Eqs. (1–2) as a function of $\gamma$ and for different matrix sizes. The probability is larger at more negative values of $\gamma$, moreover, it increases with the system size. Once concludes that Hopf bifurcations are the rule rather than the exception for sufficiently large ecosystems with $\gamma<0$.
Figure 5: Numerically computed probability that the eigenvalue with largest real part of a random matrix defined by Eq. (2) is imaginary, as a function of $\gamma$. Data correspond to averages over exact diagonalizations of 10000 matrices of size $S=57$ (violet), 157 (blue) and 557 (green), and of 5000 matrices with $S=1057$ (red) and 2057 (black). Note the limiting cases (not shown) $P(\gamma=-1)=1$ and $P(\gamma=1)=0$.
IV.2 Long-time asymptotic for different initial conditions
When $\sigma$ is smaller than the May bound, $\sigma<1/(1+\gamma)$, the long-time asymptotic is a single, globally attractive fixed point. We illustrate that, at larger $\sigma$, and with a finite (though small) extinction threshold, $N_{c}=10^{-20}$, different asymptotic behaviors can be reached. Fig. 6 shows five different dynamics obtained from five different initial distribution of populations subjected to the same interaction matrix ${\mathbb{A}}_{ij}$. For clarity, we show only five species for each case. Two initial conditions converge toward a limit cycle and three toward a fixed point. The three fixed points are evidently different. The two cycles differ mostly by the survival of species #75 in panel (a) and its extinction in panel (e). As a consequence, the oscillations are similar, but with larger amplitude in panel (e). We stress that the dynamics has been investigated for much longer times than shown, and that in all cases, the dynamics remain the same as for $t\gtrsim 350$, in particular with the same amplitudes of oscillations in panels (a) and (e).
Setting a higher extinction threshold $N_{c}=10^{-10}$ makes species #75 disappear also in panel (a), however it further modifies the dynamics so that both the cases of panel (a) and (e) converge toward different fixed points (not shown). While the specifically followed dynamics depends on $N_{c}$, the general conclusion that different dynamics can be followed by different initial conditions is general.
Figure 6: Panels (a-e): Five different dynamics, corresponding to the same realization of the interaction matrix, with five different initial distribution of populations. The limit cycles reached in panels (a) and (e) differ by the presence of one additional species in panel (a). Species oscillate about almost the same average values, with however large amplitude in panel (e). Panel (f): Dynamics of species #75 which persists in panel (a) but goes extinct in panel (e). Changing the extinction threshold to $N_{c}=10^{-10}$ leads to two different fixed points for the initial conditions of panels (a) and (e). Five species are shown for each panel. The number of surviving species is $N_{s}=57$ [panel (a)], 51 [panel (b)], 49 [panel (c)], 46 [panel (d)] and 56 [panel (e)].